OpenStudy (anonymous):

what is the area of larger segment of circle formed by chord subtending 90 degree at center and has a length of 5 cm ?

OpenStudy (anonymous):

hello

OpenStudy (primeralph):

Well, (90/360)*2pi*r = 5. r = 10/pi Area = (90/360)*pi*r^2.

OpenStudy (da_scienceman):

The chord which subtends an angle of 90 at the center of a circle might result in an isoceles triangle. Thus, the remaining larger segment of the circle area can be gotten by subtracting the area of the isoceles triangle formed by the chord. I assume the sides of this triangle will be the radius of the circle. Maybe d solution of @primeralph is correct?

OpenStudy (primeralph):

Actually my solution is incorrect.